### v2

@@ Abstract
Context: Galaxies migrate along filaments of the cosmic web from small groups to clusters, which creates the appearance that the evolution of their properties speeds up as environments get denser.
Aims: We introduce the Muse gAlaxy Groups in Cosmos (MAGIC) survey, which was built to study the impact of environment on galaxy evolution down to low stellar masses over the last 8 Gyr.
Methods: The MAGIC survey consists of 17 Multi-Unit Spectrocopic Exporer (MUSE) fields targeting 14 massive, known structures at intermediate redshift () in the COSMOS area, with a total on-source exposure of 67h. We securely measured the redshifts for 1419 sources and identified 76 galaxy pairs and 67 groups of at least three members using a friends-of-friends algorithm. The environment of galaxies is quantified from group properties, as well as from global and local density estimators.
Results: The MAGIC survey has increased the number of objects with a secure spectroscopic redshift over its footprint by a factor of about 5 compared to previous extensive spectroscopic campaigns on the COSMOS field. Most of the new redshifts have apparent magnitudes in the band . The spectroscopic redshift completeness is high: in the redshift range of [O ii] emitters (), where most of the groups are found, it globally reaches a maximum of 80% down to , and locally decreases from % to % in magnitude bins from to . We find that the fraction of quiescent galaxies increases with local density and with the time spent in groups. A morphological dichotomy is also found between bulge-dominated quiescent and disk-dominated star-forming galaxies. As environment gets denser, the peak of the stellar mass distribution shifts towards , and the fraction of galaxies with decreases significantly, even for star-forming galaxies. We also highlight peculiar features such as close groups, extended nebulae, and a gravitational arc.
Conclusions: Our results suggest that galaxies are preprocessed in groups of increasing mass before entering rich groups and clusters. We publicly release two catalogs containing the properties of galaxies and groups, respectively.
MAGIC: Muse gAlaxy Groups In Cosmos - A survey to probe the impact of environment on galaxy evolution over the last 8 Gyr ††thanks: Based on observations made with ESO telescopes at the Paranal Observatory under programs 094.A-0247, 095.A-0118, 096.A-0596, 097.A-0254, 098.A-0017, 099.A-0246, 0100.A-0607, 0101.A-0282, 0102.A-0327, 0103.A-0563. ††thanks: Catalogs described in Appendix A are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@
To make progress in our understanding of galaxy evolution, it is necessary to elucidate spatially resolved properties, such as metallicity and kinematics, inferred from spectral features. In particular, the kinematics can be used to relate galaxy integrated and morphological properties to their dynamical mass and angular momentum. Currently, the main samples for which such resolved properties are observed at intermediate to high redshifts are drawn from integral field spectrographs. Recently, the K-band multi-objet spectrograph (KMOS) has provided samples containing thousands of galaxies (KMOS, Wisnioski et al., 2015; and the KMOS redshift one spectroscopic survey, KROSS, Stott et al., 2016). However, despite their large statistics, those samples are not well suited to study the impact of environment on galaxy properties. Instead, dedicated surveys are necessary. Some have been achieved, mostly around massive clusters at intermediate redshift with KMOS (Böhm et al., 2020; Demarco et al., 2015; K-CLASH, Tiley et al., 2020; KMOS cluster survey, KCS, and the Virial survey Wilman et al., 2015; Beifiori et al., 2017; Chan et al., 2018), but also with long-slit spectrographs such as VIMOS and DEIMOS (Pelliccia et al., 2019), or FORS2 (Pérez-Martínez et al., 2017). These surveys are either focused on passive galaxies or are limited to relatively massive galaxies due to the necessary preselection of targets, similarly to the spectroscopic surveys mentioned above. Only wide-field integral field spectrographs alleviate the need for preselection. This is the case, for instance, for the study of clusters at with SITELLE at the Canada-France-Hawaii Telescope (Liu et al., 2021; Edwards et al., 2021) and of a cluster at with OSIRIS on the Gran Telescopio Canarias (Pérez-Martínez et al., 2021). In principle, such instruments can also unveil the intracluster or intragroup medium in emission. Nevertheless, these scanning instruments are of limited sensitivity. Recently, using the Multi-Unit Spectroscopic Explorer (MUSE, Bacon et al., 2010), the Middle Ages Galaxy Properties with Integral field spectroscopy (MAGPI, Foster et al., 2021) survey observed 56 fields around massive galaxies at from the GAMA survey in a range of environments. Richard et al. (2021) also observed 12 massive lensing clusters with MUSE at redshifts , some of which host ionized gas nebulae related to ram-pressure-stripping events (Moretti et al., 2022).
@@
Since MUSE has been operating on the Very Large Telescope (VLT), deep fields have proven fruitful for the discovery of groups at intermediate redshifts (, e.g., Bacon et al., 2015, 2017; Fossati et al., 2019; Leclercq et al., 2022; Johnson et al., 2022; Cherrey et al., 2023). This is thanks to the sensitivity of this instruments, and to its one square arcminute field of view (FoV), which represents from 250 to 500 kpc from to . In this redshift range, MUSE is also quite efficient in unambiguously identifying faint and low-mass galaxies (e.g., Boogaard et al., 2018; Bacon et al., 2023) and in deriving spatially resolved physical properties from the emission of their ionized gas (e.g., Contini et al., 2016; Bouché et al., 2021, 2022), but also from their stellar kinematics in the most massive cases (Guérou et al., 2017).
@@
In Sect. 2 of the present paper, we first present the MAGIC target selection, observing strategy, and data reduction, as well properties inferred from ancillary data. In Sect. 3, we explain how redshifts were determined and how MAGIC compares to previous spectroscopic campaigns, and assess its spectroscopic redshifts completeness. Section 4 is devoted to the identification of structures, to the investigation of their properties, and to the description of MAGIC group and galaxy catalogs. In Sect. 5, we present an analysis of galaxy population properties as a function of environment, including the red fraction, stellar mass, star formation rate, color, and morphology distributions. Finally, we highlight noticeable features found in MAGIC datacubes in Sect. 6, before summarizing and presenting our conclusions in Sect. 7. Throughout the paper, we assume a CDM cosmology with H km s, , and , and all physical distances mentioned in the paper are proper distances.
@@ 2.1 Target selection
Fourteen structures in the redshift range were chosen from the zCOSMOS 20k group catalog (Knobel et al., 2009, 2012)111One additional MUSE field targeting a group at from the VIMOS Very Deep Survey (VVDS) group catalog (Cucciati et al., 2010, VVDS group number 189) was also initially observed as part of this program. We decide not to include it in the MAGIC sample release for the sake of data homogeneity.. We first used this catalog to select groups containing at least five spectroscopic members within one single MUSE field, and further put emphasis on the richest groups including the largest number of emission line galaxies, with at least one emission line with a flux erg s cm. Ten groups were chosen with this strategy, among which, four groups were observed from P94 to P96, with exposures ranging from about 2 to 10 hours. After these periods, we refined our strategy to prepare for adaptive optics (AO) assisted observations: in P97 to P99, we only observed 1 hour snapshots on fields with suitable tip-tilt stars to further select the most promising candidates for subsequent AO observations. With this aim, we searched for additional fields, by allowing for mosaicking around large overdensities identified from the zCOSMOS 20k group catalog (Knobel et al., 2012) and COSMOS2015 photometric catalog (Laigle et al., 2016). This added four new fields on two main structures, including a new one (CGr32). After this snapshot campaign, we decided to discard four groups from further AO assisted observations, starting in P100. We finally decided to add three more groups from the COSMOS Wall group catalog (Iovino et al., 2016) for the last P102 and P103 periods. We selected the richest groups that could be observed with AO and that had X-ray counterparts (Gozaliasl et al., 2019)222Two of those groups would have been selected without the X-ray criterion, owing to their richness., since the densest groups we previously observed were fulfilling these criteria.
@@ 3.1 Redshifts determination
Redshifts were estimated with a MUSE customized version of the redshift-finding algorithm MARZ (Hinton et al., 2016; Inami et al., 2017) using both absorption and emission features. Redshift precision depends on (i) the nature (emission lines, absorption lines or continuum) and number of spectral features used for redshift determination, (ii) the spectral resolution at the wavelength of those features, (iii) the signal-to-noise ratio (S/N), (iv) the object internal kinematics within the aperture used for redshift determination since it may bias the systemic redshift estimate. Indeed, the absorption lines positions are biased towards the brightest regions, usually at the center of galaxies, whereas emission lines can be dominated by off-center star-forming clumps. We therefore expect a larger uncertainty for massive galaxies with asymmetric, clumpy and off-centered flux distributions in emission lines, at the exception of those seen nearly face-on. To minimize this effect, we correct the redshift of each spatially resolved galaxy for the systemic redshift derived from the best-fit model of the [O ii] kinematics performed in Mercier et al. (2023), when available. Given the typical S/N and spectral resolution of MUSE (), the redshift accuracy at 1- is about , corresponding to a 1- velocity accuracy of km s.
@@ 4 Structures identification and density estimators
The quantification of environment and the identification of structures can be done in many ways. These methods include Voronoi tesselations and Delaunay triangulation (e.g., Marinoni et al., 2002), weighted adaptive kernel smoothing, density at the radius of the n neighbor (e.g., Wang et al., 2020), friends-of-friends (FoF) algorithms, and are based either on spectroscopic or photometric redshifts or the combination of both. Each method has advantages, such as the possibility to combine both photometric and spectroscopic measurements (e.g., the Voronoi tesselations Monte-Carlo mapping -VMC- developed in Lemaux et al., 2017, 2022; Hung et al., 2020, 2021), or an easy implementation (e.g., the FoF algorithm used in Knobel et al., 2012; Iovino et al., 2016), but most of those remain sensitive to the underlying completeness of redshift measurements and need extensive sets of simulations to tune the algorithm parameters. They also depend on the size of the underlying studied field. Darvish et al. (2015) describe and compare several methods within the COSMOS field using photometric redshifts and show an overall agreement between those methods. The goal of the present paper is not to further develop these methods but to have various estimates of environment and density. The methods used to quantify environment are described hereafter.
@@ 4.1 Identification of structures using a FoF algorithm
In Knobel et al. (2012) and Iovino et al. (2016), the parameters of the algorithm were tuned with simulations to mimic their target selection based on photometric measurements, and their spectroscopic redshift success rate. Our dataset does not suffer any pre-selection and has a good and homogeneous spectroscopic redshift completeness over the whole 0.2 to 1.5 redshift range (see Sect. 3.3). In addition, the redshift uncertainty in our sample is low, km s in terms of velocity for both emission and absorption line objects (see Sect. 3.1). We therefore use a simpler scheme and run the FoF algorithm using all galaxies in the MAGIC catalog with a secure redshift (confidence flag larger than or equal to 2), without any magnitude restriction. This choice allows us to infer group membership for both high-mass and low-mass galaxies. Including the latter in the group finding algorithm also helps in refining the group parameters (Sect. 4.2) thanks to better statistics. For each source, transverse proper distances and velocity differences to all other galaxies are computed. Following Iovino et al. (2016) prescriptions, we use an iterative approach with FoF parameters depending on the number of galaxies in the groups and use maximum physical and velocity separations kpc and km s for groups of seven or more members, kpc and km s for groups of six members, kpc and km s for groups of five members, kpc and km s for groups of four members, kpc and km s for groups of three members, and kpc and km s for pairs. The velocity separations are inferred from Iovino et al. (2016) FoF parameters at , providing more stringent limits than considering higher redshifts. Friends of a source are those within the maximum transverse and longitudinal separations. All friends of a friend are then considered as members of the same structure.
@@
A refined analysis about groups statistics is provided in Sect. 4.6. Our goal here is to show how robust are the most massive groups with respect to FoF parameters and galaxies magnitude. We find 26 groups with at least eight members, all having . Using more relaxed constraints (by %) on both transverse and longitudinal separations only adds one or two galaxies for 20% of those groups. Conversely, using more stringent constraints (by %) slightly reduces the number of galaxies in some of those groups, and splits three of them in smaller sub-groups. We also run the FoF algorithm using a magnitude limit of 24.5 in the band on 3″ diameter aperture, where spectroscopic redshifts completeness in MAGIC is higher than 70% per magnitude bin. This reduces the number of groups with at least eight members to 18. Nevertheless, 21 out the 26 aforementioned groups are identified with at least five members, and the five remaining groups have either three or four members (all at ). We also do the same exercise using only galaxies brighter than , to mimic the pre-selection done in Iovino et al. (2016) and only find 13 groups with eight or more members. However, similarly to the previous test, 18 out the 26 groups found with the whole galaxy catalog are identified with at least five members within this band limited catalog, six have either three or four members (all at ), and two are not found at all (previously identified as having eight and nine members). We are therefore quite confident that our FoF algorithm finds actual groups.
@@ 4.3 Local density estimator using the VMC method
In order to avoid edge effects where the data begin to fall off, a mask is applied to each slice to mask out all pixels where the density is less than 0.1 galaxy per Mpc. The median density per slice is then calculated from the unmasked region of each slice. Since the sampling might differ with redshift due (i) to the source density falling off naturally in an apparent magnitude limited sample to higher redshifts, and (ii) to variations of spectroscopic redshifts completeness, we compute the overdensity with respect to the median density over the COSMOS field computed as at each redshift slice, rather than the density . Using the overdensity mitigates these effects to a large degree. Still, both density and overdensity are provided in our final catalog. Each galaxy in our sample is thus attributed the density and overdensity of the closest pixel in the corresponding redshift slice (middle of the slice redshift range). An example of the overdensity map at the redshift of CGr32 is provided in Fig. 8.

### v3

@@ Abstract
Context. Galaxies migrate along filaments of the cosmic web from small groups to clusters, which creates the appearance that the evolution of their properties speeds up as environments get denser.
Aims. We introduce the Muse gAlaxy Groups in Cosmos (MAGIC) survey, which was built to study the impact of environment on galaxy evolution down to low stellar masses over the last 8 Gyr.
Methods. The MAGIC survey consists of 17 Multi-Unit Spectrocopic Exporer (MUSE) fields targeting 14 massive, known structures at intermediate redshift () in the COSMOS area, with a total on-source exposure of 67h. We securely measured the redshifts for 1419 sources and identified 76 galaxy pairs and 67 groups of at least three members using a friends-of-friends algorithm. The environment of galaxies is quantified from group properties, as well as from global and local density estimators.
Results. The MAGIC survey has increased the number of objects with a secure spectroscopic redshift over its footprint by a factor of about 5 compared to previous extensive spectroscopic campaigns on the COSMOS field. Most of the new redshifts have apparent magnitudes in the band . The spectroscopic redshift completeness is high: in the redshift range of [O ii] emitters (), where most of the groups are found, it globally reaches a maximum of 80% down to , and locally decreases from % to % in magnitude bins from to . We find that the fraction of quiescent galaxies increases with local density and with the time spent in groups. A morphological dichotomy is also found between bulge-dominated quiescent and disk-dominated star-forming galaxies. As environment gets denser, the peak of the stellar mass distribution shifts towards , and the fraction of galaxies with decreases significantly, even for star-forming galaxies. We also highlight peculiar features such as close groups, extended nebulae, and a gravitational arc.
Conclusions. Our results suggest that galaxies are preprocessed in groups of increasing mass before entering rich groups and clusters. We publicly release two catalogs containing the properties of galaxies and groups, respectively.
@@
To make progress in our understanding of galaxy evolution, it is necessary to elucidate spatially resolved properties, such as metallicity and kinematics, inferred from spectral features. In particular, the kinematics can be used to relate galaxy integrated and morphological properties to their dynamical mass and angular momentum. Currently, the main samples for which such resolved properties are observed at intermediate to high redshifts are drawn from integral field spectrographs. Recently, the K-band multi-objet spectrograph (KMOS) has provided samples containing thousands of galaxies (KMOS3D, Wisnioski et al., 2015; and the KMOS redshift one spectroscopic survey, KROSS, Stott et al., 2016). However, despite their large statistics, those samples are not well suited to study the impact of environment on galaxy properties. Instead, dedicated surveys are necessary. Some have been achieved, mostly around massive clusters at intermediate redshift with KMOS (Böhm et al., 2020; Demarco et al., 2015; K-CLASH, Tiley et al., 2020; KMOS cluster survey, KCS, and the Virial survey Wilman et al., 2015; Beifiori et al., 2017; Chan et al., 2018), but also with long-slit spectrographs such as VIMOS and DEIMOS (Pelliccia et al., 2019), or FORS2 (Pérez-Martínez et al., 2017). These surveys are either focused on passive galaxies or are limited to relatively massive galaxies due to the necessary preselection of targets, similarly to the spectroscopic surveys mentioned above. Only wide-field integral field spectrographs alleviate the need for preselection. This is the case, for instance, for the study of clusters at with SITELLE at the Canada-France-Hawaii Telescope (Liu et al., 2021; Edwards et al., 2021) and of a cluster at with OSIRIS on the Gran Telescopio Canarias (Pérez-Martínez et al., 2021). In principle, such instruments can also unveil the intracluster or intragroup medium in emission. Nevertheless, these scanning instruments are of limited sensitivity. Recently, using the Multi-Unit Spectroscopic Explorer (MUSE, Bacon et al., 2010), the Middle Ages Galaxy Properties with Integral field spectroscopy (MAGPI, Foster et al., 2021) survey observed 56 fields around massive galaxies at from the GAMA survey in a range of environments. Richard et al. (2021) also observed 12 massive lensing clusters with MUSE at redshifts , some of which host ionized gas nebulae related to ram-pressure-stripping events (Moretti et al., 2022).
@@
Since MUSE has been operating on the Very Large Telescope (VLT), deep fields have proven fruitful for the discovery of groups at intermediate redshifts (, e.g., Bacon et al., 2015, 2017; Fossati et al., 2019; Leclercq et al., 2022; Johnson et al., 2022; Cherrey et al., 2024). This is thanks to the sensitivity of this instruments, and to its one square arcminute field of view (FoV), which represents from 250 to 500 kpc from to . In this redshift range, MUSE is also quite efficient in unambiguously identifying faint and low-mass galaxies (e.g., Boogaard et al., 2018; Bacon et al., 2023) and in deriving spatially resolved physical properties from the emission of their ionized gas (e.g., Contini et al., 2016; Bouché et al., 2021, 2022), but also from their stellar kinematics in the most massive cases (Guérou et al., 2017).
@@
In Sect. 2 of the present paper, we first present the MAGIC target selection, observing strategy, and data reduction, as well properties inferred from ancillary data. In Sect. 3, we explain how redshifts were determined and how MAGIC compares to previous spectroscopic campaigns, and assess its spectroscopic redshifts completeness. Section 4 is devoted to the identification of structures, to the investigation of their properties, and to the description of MAGIC group and galaxy catalogs. In Sect. 5, we present an analysis of galaxy population properties as a function of environment, including the red fraction, stellar mass, star formation rate, color, and morphology distributions. Finally, we highlight noticeable features found in MAGIC datacubes in Sect. 6, before summarizing and presenting our conclusions in Sect. 7. Throughout the paper, we assume a CDM cosmology with H km s-1, , and , and all physical distances mentioned in the paper are proper distances.
@@ 2.1 Target selection
Fourteen structures in the redshift range were chosen from the zCOSMOS 20k group catalog (Knobel et al., 2009, 2012)111One additional MUSE field targeting a group at from the VIMOS Very Deep Survey (VVDS) group catalog (Cucciati et al., 2010, VVDS group number 189) was also initially observed as part of this program. We decide not to include it in the MAGIC sample release for the sake of data homogeneity.. We first used this catalog to select groups containing at least five spectroscopic members within one single MUSE field, and further put emphasis on the richest groups including the largest number of emission line galaxies, with at least one emission line with a flux erg s-1 cm-2. Ten groups were chosen with this strategy, among which, four groups were observed from P94 to P96, with exposures ranging from about 2 to 10 hours. After these periods, we refined our strategy to prepare for adaptive optics (AO) assisted observations: in P97 to P99, we only observed 1 hour snapshots on fields with suitable tip-tilt stars to further select the most promising candidates for subsequent AO observations. With this aim, we searched for additional fields, by allowing for mosaicking around large overdensities identified from the zCOSMOS 20k group catalog (Knobel et al., 2012) and COSMOS2015 photometric catalog (Laigle et al., 2016). This added four new fields on two main structures, including a new one (CGr32). After this snapshot campaign, we decided to discard four groups from further AO assisted observations, starting in P100. We finally decided to add three more groups from the COSMOS Wall group catalog (Iovino et al., 2016) for the last P102 and P103 periods. We selected the richest groups that could be observed with AO and that had X-ray counterparts (Gozaliasl et al., 2019)222Two of those groups would have been selected without the X-ray criterion, owing to their richness., since the densest groups we previously observed were fulfilling these criteria.
@@ 3.1 Redshifts determination
Redshifts were estimated with a MUSE customized version of the redshift-finding algorithm MARZ (Hinton et al., 2016; Inami et al., 2017) using both absorption and emission features. Redshift precision depends on (i) the nature (emission lines, absorption lines or continuum) and number of spectral features used for redshift determination, (ii) the spectral resolution at the wavelength of those features, (iii) the signal-to-noise ratio (S/N), (iv) the object internal kinematics within the aperture used for redshift determination since it may bias the systemic redshift estimate. Indeed, the absorption lines positions are biased towards the brightest regions, usually at the center of galaxies, whereas emission lines can be dominated by off-center star-forming clumps. We therefore expect a larger uncertainty for massive galaxies with asymmetric, clumpy and off-centered flux distributions in emission lines, at the exception of those seen nearly face-on. To minimize this effect, we correct the redshift of each spatially resolved galaxy for the systemic redshift derived from the best-fit model of the [O ii] kinematics performed in Mercier et al. (2023), when available. Given the typical S/N and spectral resolution of MUSE (), the redshift accuracy at 1- is about , corresponding to a 1- velocity accuracy of km s-1.
@@ 4 Structures identification and density estimators
The quantification of environment and the identification of structures can be done in many ways. These methods include Voronoi tesselations and Delaunay triangulation (e.g., Marinoni et al., 2002), weighted adaptive kernel smoothing, density at the radius of the nth neighbor (e.g., Wang et al., 2020), friends-of-friends (FoF) algorithms, and are based either on spectroscopic or photometric redshifts or the combination of both. Each method has advantages, such as the possibility to combine both photometric and spectroscopic measurements (e.g., the Voronoi tesselations Monte-Carlo mapping -VMC- developed in Lemaux et al., 2017, 2022; Hung et al., 2020, 2021), or an easy implementation (e.g., the FoF algorithm used in Knobel et al., 2012; Iovino et al., 2016), but most of those remain sensitive to the underlying completeness of redshift measurements and need extensive sets of simulations to tune the algorithm parameters. They also depend on the size of the underlying studied field. Darvish et al. (2015) describe and compare several methods within the COSMOS field using photometric redshifts and show an overall agreement between those methods. The goal of the present paper is not to further develop these methods but to have various estimates of environment and density. The methods used to quantify environment are described hereafter.
@@ 4.1 Identification of structures using a FoF algorithm
In Knobel et al. (2012) and Iovino et al. (2016), the parameters of the algorithm were tuned with simulations to mimic their target selection based on photometric measurements, and their spectroscopic redshift success rate. Our dataset does not suffer any pre-selection and has a good and homogeneous spectroscopic redshift completeness over the whole 0.2 to 1.5 redshift range (see Sect. 3.3). In addition, the redshift uncertainty in our sample is low, km s-1 in terms of velocity for both emission and absorption line objects (see Sect. 3.1). We therefore use a simpler scheme and run the FoF algorithm using all galaxies in the MAGIC catalog with a secure redshift (confidence flag larger than or equal to 2), without any magnitude restriction. This choice allows us to infer group membership for both high-mass and low-mass galaxies. Including the latter in the group finding algorithm also helps in refining the group parameters (Sect. 4.2) thanks to better statistics. For each source, transverse proper distances and velocity differences to all other galaxies are computed. Following Iovino et al. (2016) prescriptions, we use an iterative approach with FoF parameters depending on the number of galaxies in the groups and use maximum physical and velocity separations kpc and km s-1 for groups of seven or more members, kpc and km s-1 for groups of six members, kpc and km s-1 for groups of five members, kpc and km s-1 for groups of four members, kpc and km s-1 for groups of three members, and kpc and km s-1 for pairs. The velocity separations are inferred from Iovino et al. (2016) FoF parameters at , providing more stringent limits than considering higher redshifts. Friends of a source are those within the maximum transverse and longitudinal separations. All friends of a friend are then considered as members of the same structure.
@@
A refined analysis about groups statistics is provided in Sect. 4.6. Our goal here is to show how robust are the most massive groups with respect to FoF parameters and galaxies magnitude. We find 26 groups with at least eight members, all having . Using more relaxed constraints (by %) on both transverse and longitudinal separations only adds one or two galaxies for 20% of those groups. Conversely, using more stringent constraints (by %) slightly reduces the number of galaxies in some of those groups, and splits three of them in smaller sub-groups. We also run the FoF algorithm using a magnitude limit of 24.5 in the band on 3″ diameter apertures, where spectroscopic redshifts completeness in MAGIC is higher than 70% per magnitude bin. This reduces the number of groups with at least eight members to 18. Nevertheless, 21 out the 26 aforementioned groups are identified with at least five members, and the five remaining groups have either three or four members (all at ). We also do the same exercise using only galaxies brighter than , to mimic the pre-selection done in Iovino et al. (2016) and only find 13 groups with eight or more members. However, similarly to the previous test, 18 out the 26 groups found with the whole galaxy catalog are identified with at least five members within this band limited catalog, six have either three or four members (all at ), and two are not found at all (previously identified as having eight and nine members). We are therefore quite confident that our FoF algorithm finds actual groups.
@@ 4.3 Local density estimator using the VMC method
In order to avoid edge effects where the data begin to fall off, a mask is applied to each slice to mask out all pixels where the density is less than 0.1 galaxy per Mpc2. The median density per slice is then calculated from the unmasked region of each slice. Since the sampling might differ with redshift due (i) to the source density falling off naturally in an apparent magnitude limited sample to higher redshifts, and (ii) to variations of spectroscopic redshifts completeness, we compute the overdensity with respect to the median density over the COSMOS field computed as at each redshift slice, rather than the density . Using the overdensity mitigates these effects to a large degree. Still, both density and overdensity are provided in our final catalog. Each galaxy in our sample is thus attributed the density and overdensity of the closest pixel in the corresponding redshift slice (middle of the slice redshift range). An example of the overdensity map at the redshift of CGr32 is provided in Fig. 8.